OpenStudy (anonymous):

8^(log(base2)x)-25^(log(base5)x)=4x-4

OpenStudy (anonymous):

mmm that is good question.

OpenStudy (anonymous):

(2^3) ^(log2 (x) ) - (5^2) ^(log5 (x))= 4x -4

OpenStudy (anonymous):

2^(log2 (x^3) ) - 5^(log5 (x^2) ) = 4x-4

OpenStudy (anonymous):

x^3-x^2-4x+4=0 x^2(x-1)-4(x-1)=0 (x-1)(x-2)(x+2)=0 so x=1,2,-2

OpenStudy (anonymous):

but, check your answers in the eqn

OpenStudy (anonymous):

you cannot take log of a negative. So your answers are x=1,2

OpenStudy (anonymous):

hopefully you follow the working out. Its just basic algebra

OpenStudy (anonymous):

it works. thanx for the help

OpenStudy (anonymous):

its just thinks like a^(log a *(x) ) = x [ logorithms and indices are inverses] (x^a)^b = x^(ab) and n log (x) = log (x^n) and seeing the patterns

OpenStudy (anonymous):

It is one of the hardest high school questions I have seen on logs though.

OpenStudy (anonymous):

your smart. thanks for helping me out. :)))